# Detection of Deepfake Environmental Audio

### v1

@@ Abstract
With the ever-rising quality of deep generative models, it is increasingly important to be able to discern whether the audio data at hand have been recorded or synthesized. Although the detection of fake speech signals has been studied extensively, this is not the case for the detection of fake environmental audio.
We propose a simple and efficient pipeline for detecting fake environmental sounds based on the CLAP audio embedding. We evaluate this detector using audio data from the 2023 DCASE challenge task on Foley sound synthesis.
Our experiments show that fake sounds generated by 44 state-of-the-art synthesizers can be detected on average with 98% accuracy. We show that using an audio embedding learned on environmental audio is beneficial over a standard VGGish one as it provides a 10 % increase in detection performance. Informal listening to Incorrect Negative examples demonstrates audible features of fake sounds missed by the detector such as distortion and implausible background noise.
@@ I Introduction
While deepfake detection for speech is well studied, there appears to be little research, to our knowledge, on the detection of fake content for environmental sounds. This paper aims to fill this gap with respect to Foley sound synthesis, i.e., sound categories such as hand clap, rain, etc.
@@
The paper is organized as follows. Section II gives a brief review of the relevant state of the art on deepfake detection in audio. Section III introduces the proposed deepfake detector which is benchmarked using an experimental protocol described in Section IV. Performance is discussed in Section V and Section VI discusses the outcomes of informal listening of detection mistakes of the proposed detector that suggest directions for improvement. Code, supplementary material, and audio examples are available on the companion page  [^1] Companion page: https://mathieulagrange.github.io/audioFoleyDeepFake.
@@ II Related Work
This section discusses machine-learning techniques for Deepfake Audio detection. In [10], the authors use linear regression to detect fake voices. The features are extracted using a signal’s entropy calculated based on Shannon’s equation for each second of the signal and for the whole signal. The model reaches 98% accuracy, with all fake audio being correctly detected. In [11], the authors adopt a combination of Bispectral analysis and MFCC (Mel Frequency Cepstral Coefficients) for speech detection and feature extractions. The former helps to identify components in generated audio that are not present in human speech and the latter is used to detect the features in human speech tied to the vocal tract that are absent in the AI-generated speech. The features are then fed into seven different ML algorithms. The results show that the Quadratic SVM model gives the highest accuracy of 96.1%. In [12], the feature vector is extracted by filtering the signal under analysis and extracting different statistics with short-term and long-term prediction order. The features are then faded to a Linear SVM and Random Forest model. The results show that the SVM outperforms the RF in different conditions.
@@ III Proposed approach
The detection of fake sounds is treated as a binary classification task. As can be seen in Figure 1, the proposed architecture for solving this task leverages the power of pre-trained audio embeddings. This allows us to simplify the learning process by requiring less training data and using less power as the learnable part of the detector consists of only a few Multilayer Perceptron (MLP) or Convolutional Neural Network (CNN) layers.
The CNN architecture consists of two sequential convolutional layers with ReLU activations and subsequent max pooling for feature abstraction, with a flattening layer before a terminal linear transformation. However, due to the poor performance of CNNs, we focus exclusively on reporting the performance of the MLP model. This model is structured, as shown in 1, with three dense layers. Each linear layer is followed by 1) ReLu activation functions to reduce the backpropagation errors and accelerate the learning process, and 2) dropout layers for regularization to avoid overfitting. A final dense layer with a sigmoid function as the output activation function.
@@ IV-A Dataset
The resulting dataset consists of both nonfake and fake sounds across seven distinct sound classes (dog_bark, footstep, gunshot, keyboard, moving_motor_vehicle, rain, and sneeze_cough). Specifically, the dataset comprises 5,550 nonfake sounds alongside 25,200 fake sounds, the latter of which were generated by challenge participants. These sounds are analyzed through the lens of four different embeddings, detailed in section II.B, each with its unique dimensionality: VGGish embeddings at (4,128), Microsoft CLAP (MS-Clap) embeddings at (4,1024), and two PANNs embeddings at (4,2048), where ’4’ refers to the four-second duration of the audio clips. For training and evaluation purposes, we separate the dataset into a training set (70%), a validation set (10%), and an evaluation set (20%)  [^2] Dataset of recorded and generated audio: https://zenodo.org/records/8091972.
MLP VGGish MLP MS-Clap MLP PANN_Wavegram_logmel MLP PANN_32k
@@
TABLE I: Confusion Matrix Evaluation for MLP models
@@ IV-B Training Procedure
A simple preprocessing step is performed on the embeddings: we take the mean (by axis 0) of each embedding to avoid the temporal sequence aspect (four-second duration) affecting the deepfake detection. Figure 1 shows the pipeline used in the experiment protocol. As we have four different input embeddings, the procedure yielded four models to use and compare: MLP VGGish, MLP MS-Clap, MLP PANN Wavegram, and MLP PANNcnn14-32k. These models were trained using the Adam optimizer, configured with a learning rate of 7e-4. The training protocol was standardized across models, with a batch size of 128 samples, a decision threshold of 0.5 on both training and validation predictions, and training over 100 epochs. Throughout this process, the Binary Cross-Entropy (BCE) 1 loss function is used to optimize the weights parameters in the backpropagation process:
@@ V-A Inference Time
To evaluate the computational efficiency of each model, we assessed the time required for the embedding process and the MLP model’s inference for each embedding on one random audio sample across 100 runs. The inference time (in seconds) is then computed by averaging over the 100 inference times. In Table II, the averaged inference time is expressed as a percentage of the real-time duration of the audio sample. This metric provides a clear understanding of the model’s speed relative to the actual audio duration.
@@ V-B Overall Accuracy
We calculate both the relative overall validation and evaluation accuracy, along with the time per epoch in seconds (s), and generate a Confusion Matrix for the evaluation dataset. As presented in Table II, statistics such as Overall Evaluation Accuracy and Training Time per Epoch are detailed for each model. Among the experiments conducted under the same conditions, the MLP MS-Clap model outperformed others with the highest Evaluation Accuracy of 98.02%, surpassing both MLP PANN models by 5% and the MLP VGGish model by 10%.
@@
MLP VGGish 88.11 $\pm$ 0.73 0.423
MLP MS-Clap 98.02 $\pm$ 0.18 1.82
MLP PANN-Wavegram 93.15 $\pm$ 0.34 0.318
MLP PANNcnn14_32k 93.04 $\pm$ 0.32 0.234
Among the 40 identical runs, we pick the best-performing run for each model (i.e., the highest evaluation accuracy). The confusion matrices resulting from the four chosen MLP models demonstrate a varied performance across the different architectures. The MLP VGGish model shows difficulty with correctly classifying nonfake sounds, as it produces 9% false positive judgements. The MLP MS-Clap model shows the best proficiency in identifying the nonfake category of sounds, with the lowest false prediction rate (2% of the dataset). On the other hand, the MLP PANN-Wavegram and the MLP PANN_32k models behave approximately the same by having a high evaluation accuracy and also a low false prediction rate, yet they do not perform as well overall as the MLP Clap model.
@@ V-C Statistical Analysis
The accuracy of MLP MS-Clap is higher than that of MLP Pann-Wavegram with 5.02%, and that of MLP PANN 32k with 5.11%, and that of MLP VGGish with 10.23%. In all cases, the U-test confirms that those differences are significant, as for all pairs the Null hypothesis is rejected.
@@
TABLE III: Per Class Accuracy of the MLP MS-Clap deepfake predictor.
@@ V-D Class-Wise Accuracy
As the MLP MS-Clap model is the best-performing model among the four trained models, we chose to carry out our analysis with this model. To have a better understanding of the behavior of the model’s performance, we look closely at the model’s accuracy for each class (Table III). This closer look helps us understand how the model performed with different types of data. We find that the model was consistently good at identifying all classes (98% approximately for each class), with dog_bark being the class where the model performs the best, and it performs slightly less better than the others when it has to predict on keyboard and sneeze_cough classes. [^3] On the companion website a confusion matrix is shown for each evaluated embedding. Entries of those matrices are per system generator and sound class. This matrix shows that it is challenging to detect some fake sounds that are generated by some specific systems on specific classes.
@@
VI Discussion
We now discuss the outcome of informal listening tests that were performed to qualitatively evaluate the acoustic properties of A) incorrect positives, and B) incorrect negatives.
VI-A Incorrect Positives
@@
Erroneously classifying a nonfake sound as a fake is an ”Incorrect Positive” (or ”False alarm”) whereas classifying a fake sound as a fake is a ”Correct Positive” (or ”Hit”). We examined 38 Incorrect Positive sounds, spanning all 7 categories, based on the erroneous high likelihood of the MLP MS-Clap model. These sounds are from the DCASE development and evaluation dataset that were presumed to be real recordings; however, although these sounds were aurally screened by the organizers, the datasets may have nonetheless contained a few Foley or processed sounds. To detect this possibility and also to detect other reasons the system incorrectly classified a sound as fake, all authors carefully listened to these sounds. We noted two recordings that had very strong echoes. We suspected that two recordings were Foley-simulated footsteps rather than recordings of real footsteps. Some recordings included non-target sounds (e.g. footsteps during rain, or a snore after a sneeze). Many sounds had significant background noise. We verbally characterized five sound features that may have triggered an incorrect fake detection and used them to create a menu of ”reasons” for a person suspecting that an audio is fake. Next, one naive lab member blindly rated all 38 sounds as real or fake (presented in random order along with the Incorrect Negative samples). A response of ”fake” brought up a menu of five possible reasons for their judgement: I hear extra artifacts that do not belong; The sound is unrealistic; The sound does not fit the category; The sound is too repetitive; The audio does not sound like it is from a recording; Other. The most common reason given was unrealistic sound, followed by audible artifacts, followed by not fitting the category. No other reasons were selected. The listener correctly recognized 63.2% of these Incorrect Positives as nonfake, which was marginally greater than a chance rate of 50% (p$<$.07, Binomial test) and significantly greater than the 0% accuracy of our fake detector (p$<$.00001).
VI-B Incorrect Negatives
@@
The 43 fake sounds that were erroneously classified as nonfake by the MLP MS-Clap detector are referred to as ”Incorrect Negatives”. There were sounds from all 7 categories. We noted a couple of recordings that contained very little sound (perhaps a brief thud) and so there was not much available acoustic information. Many sounds had significant background noise, and often the noise sounded more like static than the typical noise heard in real recordings, such as environmental noise or distortion from overloading a microphone (clipping/saturation). Sometimes the onset of the target sound contained a burst of noise. One sound had audible repetition; evidence of splicing/editing was confirmed via a spectrogram. Next, as described above, one lab member blindly rated all Incorrect Negatives as real or fake (presented in random order along with the Incorrect Positives). The most common reason given was audible artifacts, followed by unrealistic sound, followed by failure to fit the category, followed by too repetitive. The overall pattern of reasons for a fake judgement was similar for the Incorrect Negatives and Incorrect Positives. The listener correctly recognized 81.4% of Incorrect Negatives as fake, which was significantly greater than a chance rate of 50% (p$<$.00004, Binomial test) and significantly greater than the 0% accuracy of our fake detector (p$<$.00001). Their overall accuracy was 72.8%. Using the signal detection theory framework, d’ is a measure of sensitivity to fake sounds which is independent of the bias to decide real/fake. By design, our detector had zero sensitivity to this set of 81 sounds, whereas our naive listener displayed notable sensitivity with a d’=1.23 (derived from a hit rate = 0.814 and a false alarm rate = 0.368, i.e. the rate at which the listener made errors on the set of Incorrect Positives.) In conclusion, we provide a ”proof of concept” that there are audible acoustic cues that help distinguish fake and nonfake sounds which have not yet been exploited by our deepfake detection system. Human listening suggests opportunities for improvement in the detection of distortion, realistic echoes and temporal structure, and patterns of noise and repetition which may be noticeable to humans but not to a frame-based detector such as the one considered in this study.
VII Acknowledgments
@@
The authors thank Justine Sullivan for helping with the blind listening test described in Section VI.

### v2

@@ Abstract
With the ever-rising quality of deep generative models, it is increasingly important to be able to discern whether the audio data at hand have been recorded or synthesized. Although the detection of fake speech signals has been studied extensively, this is not the case for the detection of fake environmental audio. We propose a simple and efficient pipeline for detecting fake environmental sounds based on the CLAP audio embedding. We evaluate this detector using audio data from the 2023 DCASE challenge task on Foley sound synthesis.
Our experiments show that fake sounds generated by 44 state-of-the-art synthesizers can be detected on average with 98% accuracy. We show that using an audio embedding trained specifically on environmental audio is beneficial over a standard VGGish one as it provides a 10% increase in detection performance. The sounds misclassified by the detector were tested in an experiment on human listeners who showed modest accuracy with nonfake sounds, suggesting there may be unexploited audible features.
@@ I Introduction
While deepfake detection for speech is well studied, there appears to be little research on the detection of fake content for environmental sounds. This paper aims to fill this gap with respect to Foley sound synthesis, i.e., sound categories such as hand clap, rain, etc.
@@
The paper is organized as follows. Section II gives a brief review of the relevant state of the art on deepfake detection in audio. Section III introduces the proposed deepfake detector which is benchmarked using an experimental protocol described in Section IV. Performance is discussed in Section V and Section VI discusses the outcome of listening tests on detection mistakes of the proposed detector and directions for improvement. Code, supplementary material, and audio examples are available on the companion page. [^1] Companion page: https://mathieulagrange.github.io/audioFoleyDeepFake
@@ II Related Work
This section discusses machine-learning techniques for Deepfake Audio detection. In [10], the authors use linear regression to detect fake voices. The features are extracted using a signal’s entropy calculated based on Shannon’s equation for each second of the signal and for the whole signal. The model reaches 98% accuracy, with all fake audio being correctly detected. In [11], the authors adopt a combination of Bispectral analysis and MFCC (Mel Frequency Cepstral Coefficients) for speech detection and feature extractions. The former helps to identify components in generated audio that are not present in human speech and the latter is used to detect the features in human speech tied to the vocal tract that are absent in the AI-generated speech. The features are then fed into seven different ML algorithms. The results show that the Quadratic SVM model gives the highest accuracy of 96.1%. In [12], the feature vector is extracted by filtering the signal under analysis and extracting different statistics with short-term and long-term prediction order. The features are then fed to a Linear SVM and Random Forest model. The results show that the SVM outperforms the RF in different conditions.
@@ III Proposed approach
The detection of fake sounds is treated as a binary classification task. As can be seen in Figure 1, the proposed architecture for solving this task leverages the power of pre-trained audio embeddings. This allows us to simplify the learning process by requiring less training data and using less power as the learnable part of the detector consists of three dense layers, as shown in Section 1. Each layer is followed by 1) ReLU activation functions to reduce the backpropagation errors and accelerate the learning process, and 2) dropout layers for regularization to avoid overfitting. The final dense layer is followed by a sigmoid function as the output activation function.
@@ IV-A Dataset
The resulting dataset consists of both nonfake and fake sounds across seven distinct sound classes (dog_bark, footstep, gunshot, keyboard, moving_motor_vehicle, rain, and sneeze_cough). Specifically, the dataset comprises 5,550 nonfake sounds alongside 25,200 fake sounds, the latter of which were generated by challenge participants. These sounds are analyzed through the lens of four different embeddings, detailed in section II.B, each with its specific dimensionality: VGGish embeddings at (4,128), Microsoft CLAP (MS-Clap) embeddings at (4,1024), and two PANNs embeddings at (4,2048), where ’4’ refers to the four-second duration of the audio clips. For training and evaluation purposes, we separate the dataset into a training set (70%), a validation set (10%), and an evaluation set (20%)  [^2] Dataset of recorded and generated audio: https://zenodo.org/records/8091972.
VGGish MS-Clap PANN_Wavegram_logmel PANN_32k
@@
TABLE I: Confusion Matrix Evaluation for the three embeddings
@@ IV-B Training Procedure
Figure 1 shows the pipeline used in the experiment protocol. As we have four different input embeddings, the procedure yielded four models to use and compare: VGGish, MS-Clap, PANN Wavegram, and PANNcnn14-32k. Note that we take the time average of each embedding to avoid the temporal sequence aspect (four-second duration) affecting the deepfake detection.
These models were trained using the Adam optimizer, configured with a learning rate of 7e-4. The training protocol was standardized across models, with a batch size of 128 samples, a decision threshold of 0.5 on both training and validation predictions, and training over 100 epochs. Throughout this process, the Binary Cross-Entropy (BCE) loss function given in Equation 1 is used to optimize the weights parameters in the backpropagation process:
@@ V-A Inference Time
To evaluate the computational efficiency of each model, we assessed the time required for the embedding process and the model’s inference for each embedding on one random audio sample across 100 runs. The inference time (in seconds) is then computed by averaging over the 100 inference times. In Table II, the averaged inference time is expressed as a percentage of the real-time duration of the audio sample. This metric provides a clear understanding of the model’s speed relative to the actual audio duration.
@@ V-B Overall Accuracy
We calculate both the relative overall validation and evaluation accuracy, along with the time per epoch in seconds (s), and generate a Confusion Matrix for the evaluation dataset. As presented in Table II, statistics such as Overall Evaluation Accuracy and Training Time per Epoch are detailed for each model. Among the experiments conducted under the same conditions, the MS-Clap model outperformed others with the highest Evaluation Accuracy of 98.02%, surpassing both PANN models by 5% and the VGGish model by 10%.
@@
VGGish 88.11 $\pm$ 0.73 0.423
MS-Clap 98.02 $\pm$ 0.18 1.82
PANN-Wavegram 93.15 $\pm$ 0.34 0.318
PANNcnn14_32k 93.04 $\pm$ 0.32 0.234
Among the 40 identical runs, we pick the best-performing run for each model (i.e., the highest evaluation accuracy). The confusion matrices resulting from the four chosen models demonstrate a varied performance across the different architectures. The VGGish model shows difficulty with correctly classifying nonfake sounds, as it produces 9% false positive judgements. The MS-Clap model shows the best proficiency in identifying the nonfake category of sounds, with the lowest false prediction rate (2% of the dataset). On the other hand, the PANN-Wavegram and the PANN_32k models behave approximately the same by having a high evaluation accuracy and also a low false prediction rate, yet they do not perform as well overall as the Clap model.
@@ V-C Statistical Analysis
The accuracy of MS-Clap is higher than that of Pann-Wavegram with 5.02%, and that of PANN 32k with 5.11%, and that of VGGish with 10.23%. In all cases, the U-test confirms that those differences are significant, as for all pairs the Null hypothesis is rejected.
@@
TABLE III: Per Class Accuracy of the MS-Clap deepfake predictor.
@@ V-D Class-Wise Accuracy
As the MS-Clap model is the best-performing model among the four trained models, we chose to carry out our next analysis with this model. To have a better understanding of the behavior of the model’s performance, we look closely at the model’s accuracy for each class (Table III). This closer look helps us understand how the model performed with different types of data. We find that the model was consistently good at identifying all classes (98% approximately for each class), with dog_bark being the class where the model performs the best, and it performs slightly worse than the others when it has to predict on keyboard and sneeze_cough classes.
@@
VI Listening Test
Human listening suggests opportunities for improvement in the detection of distortion, realistic echoes and temporal structure, and patterns of noise and repetition which may be noticeable to humans but not to a frame-based detector such as the one considered in this study.
VI-A Data
Incorrect Positives
@@ VI-A Data
Erroneously classifying a nonfake sound as a fake is an ”Incorrect Positive” whereas classifying a fake sound as a fake is a ”Correct Positive”. We examined 38 Incorrect Positive sounds, spanning all 7 categories, based on the erroneous high likelihood of the MS-Clap model. These sounds are from the DCASE development and evaluation dataset that were presumed to be real recordings; however, although these sounds were aurally screened by the organizers, the datasets may have nonetheless contained a few Foley or processed sounds. To detect this possibility and also to detect other reasons the system incorrectly classified a sound as fake, all authors carefully listened to these sounds. We noted two recordings that had very strong echoes. We suspected that two recordings were Foley-simulated footsteps rather than recordings of real footsteps. Some recordings included non-target sounds (e.g. footsteps during rain, or a snore after a sneeze). Many sounds had significant background noise. We verbally characterized five sound features that may have triggered an incorrect fake detection and used them to create a menu of ”reasons” for a person suspecting that an audio is fake.
Incorrect Negatives
@@
The 43 fake sounds that were erroneously classified as nonfake by the MS-Clap detector are referred to as ”Incorrect Negatives”. There were sounds from all 7 categories. We noted a couple of recordings that contained very little sound (perhaps a brief thud) and so there was not much available acoustic information. Many sounds had significant background noise, and often the noise sounded more like static than the typical noise heard in real recordings, such as environmental noise or distortion from overloading a microphone (clipping/saturation). Sometimes the onset of the target sound contained a burst of noise. One sound had audible repetition; evidence of splicing/editing was confirmed via a spectrogram.
VI-B Methods
Twenty CMU undergraduates participated for academic credit (Mean age=19.8; 9 female, 10 male, 1 nonbinary) who reported normal hearing and passed a binaural screening survey to ensure careful listening over headphones. They used an online platform (gorilla.sc) and gave online consent approved by CMU’s IRB. Each sound was accompanied by its category name on the screen (e.g. “dog bark”), and each sound was played once before a judgement was made. All 7 sound categories from the System’s test set were represented in the set of 43 fake and 38 real sounds on which the ML system made errors. Participants were instructed to respond “real” if the sound seemed like a real recording of an event, and to respond “fake” if it seemed like it was generated by a computer program. They were told that the existence of background noise was not an indicator of fakeness and that cues could be present in either the sound from the category or the background noises.
VI-C Outcomes
When a fake (model-generated) sound was missed by the ML system (a system incorrect negative), the percent of sounds that listeners judged as fake (listener correct positives, M=49%) was not significantly greater than chance, z=-0.07, p=0.6. When a nonfake (recorded) sound was judged as fake by the ML system (system incorrect positive), the percent of sounds that listeners judged as real (listener correct negatives, M=71%) was greater than chance, z=1.34, p $<$ 0.0001. This human sensitivity advantage implies that some acoustic features which indicated fakeness to the ML system did not seem artificial to human listeners. When a sound was judged as fake by a participant, they indicated one of six possible reasons in the following proportions: The audio of the target sound is distorted (underwater, wobbly): 34.2% The sound does not sound like an example of the target category. 32.2% I hear suspicious artifacts in the background (whining, etc.): 12.9%. The sound is too brief or quiet to be able to give a reason. 10.0% The sound is too repetitive. 6.8% Other: 4%.
VII Conclusion
We documented the first study of deep fake detection for environmental audio generation. The experiments reported considered a public dataset of sounds generated by 10 state of the art systems and another public dataset of environmental audio recordings.
The excellent results of the proposed deep fake detector demonstrate that, to this day, generated sounds can be detected rather easily. The listening test provide some interesting cues that can suggest research directions for the improvement of the generators.
VIII Acknowledgments
@@
The authors thank Justine Sullivan and Anjelica Ferguson for helping with the blind listening test described in Section VI.